OpenStudy (anonymous):

Please help! Medal will be given. A parabola contains the points (1,8), (4, 29), and (-3, 8). What is the equation of this parabola in standard form?

OpenStudy (jack1):

how to guide is here: http://jwilson.coe.uga.edu/emt668/emat6680.f99/jones/instructional%20unit/writingquads.html video version is here:(probablyfaster if you already understand the maths) https://www.youtube.com/watch?v=7KC8KqljahE lemme know if you have any trouble with these man, or post the answer you get and we'll compare if you like?

OpenStudy (anonymous):

@Jack1 - Thank you so much! But I got the systems of equations and I keep getting the wrong answer. I don't think I'm doing it right. Sorry.

OpenStudy (jack1):

what's the answer you keep getting if you dont mind me asking?

OpenStudy (anonymous):

I got y = 0x^2 + 7x + 1....

OpenStudy (jack1):

ok, i got y = x^2 + 2x + 5 if you post your step by step i could have a look for you?

OpenStudy (anonymous):

8 = a + b + c 29 = 16a + 4b + c 8 = 9a - 4b + c ----------------------- 8 = a + b + c -c + 8 = a + b -c = -8 + a + b c = 8 - a - b ------------------------ 29 = 16a + 4b + 8 - a - b 29 = 15a +3b + 8 21 = 15a + 3b ------------------------ 8 = 9a - 4b + 8 - a - b 8 = 8a - 3b + 8 0 = 8a - 3b ------------------------- 21 = 15a + 3b -3b = -21 + 15a b = 7 + 15a (I made an error. It should be b = 7 - 5a) ------------------------- a = 0 b = 7 c = 1

OpenStudy (anonymous):

@Jack1

OpenStudy (anonymous):

At the top, your third equation 8=9a-4b+c is incorrect. It should be 8=9a-3b+c